COMPARISON OF He AND H_{2} PRESSURE BROADENING OF NH_{3} FROM 15 TO 40 K

\underline{D. R. WILLEY}, R. E. TIMLIN, M. DERAMO, P. L. PONDILLO, D. M. WESOLEK AND R. W. WIG, Department of Physics, Allegheny College, Meadville, PA 16335.

Pressure broadening of the (J, K) = (1, 1), (2, 2) and (3, 3) inversion transitions of NH_{3} was measured using normal- H_{2} as the broadening agent at kinetic temperatures of 15 to 40 K. Measurements were taken in a quasi equilibrium cell using the collisional cooling technique. H_{2} pressure broadening cross sections were compared to low temperature He pressure broadening of the same transitions and found to be from 2.5 to 8 times larger than corresponding He cross sections. Measured normal H_{2} and He cross sections were also compared to calculated J = 0, para- H_{2} cross sections. Preliminary experimental results for broadening by para- H_{2} will also be presented.